A right triangle with sides 6cm,8cm and 10cm is revolved about the side 8cm.
Find the volume and curved surface area of solid so formed.

Grade:11

4 Answers

Vasantha Kumari
10 years ago

The given triangle is revolved around the side 8 cm and hence the radius of our new cone(r) = 6cm and height (h) = 8 cm.

Volume of cone=301.71

Curved surface Area=prl

Thanks & Regards,

Vasantha Sivaraj,

askIITians faculty.

Vasantha Kumari
10 years ago
ANSW attached
Vasantha Kumari
10 years ago
answer attached
Arpit Kumar
10 years ago
The solid so formed will be a right circular cone of height 8cm,radius 6cm,slant height 10cm. Volume=1/3*22/7*36*8=301.7cubic cm CSA=22/7*6*10=188.57square cm